OpenStudy (loser66):

Prove that \((p-3)! \equiv \dfrac{p-1}{2} mod p\) Please, help

OpenStudy (loser66):

p is prime, p >2

OpenStudy (loser66):

By Wilson, we have (p-1)! =-1 (mod p) hence (p-1)(p-2) (p-3)! = -1 (mod p) and p-1 =1 mod p p-2 = 2 mod p hence 2(p-3)! =-1 mod p then I am stuck.

jimthompson5910 (jim_thompson5910):

it should be p-1 = -1 (mod p) p-2 = -2 (mod p)

OpenStudy (loser66):

yes,

OpenStudy (loser66):

Fortunately, the result is the same and I stuck still!! :)

jimthompson5910 (jim_thompson5910):

(p-3)! = (p-1)/2 (mod p) is equivalent to 2(p-3)! = p-1 (mod p)

OpenStudy (loser66):

What if I * (-1) both sides and get -2 (p-3)! = 1 mod p that is (p-3)! is inverse of 2 mod p? then?

jimthompson5910 (jim_thompson5910):

then you can use wilson's theorem to get (p-1)! =-1 (mod p) (p-1)(p-2) (p-3)! = -1 (mod p) (-1)(-2)(p-3)! = -1 (mod p) 2(p-3)! = -1 (mod p)

OpenStudy (loser66):

@jim_thompson5910 Yes, that was what I did above.

OpenStudy (loser66):

but how to go further?

jimthompson5910 (jim_thompson5910):

well I think that's enough to show that they are equivalent `(p-3)! = (p-1)/2 (mod p)` is equivalent to `2(p-3)! = p-1 (mod p)` `(p-1)! =-1 (mod p)` turns into `2(p-3)! = -1 (mod p)` by the transitive property of equivalence, we know `(p-3)! = (p-1)/2 (mod p)` is equivalent to `(p-1)! =-1 (mod p)` so because `p-1)! =-1 (mod p)` is true, this makes `(p-3)! = (p-1)/2 (mod p)` true too

OpenStudy (loser66):

aah it drove me crazy :)

OpenStudy (loser66):

first: we don't have (p-3)! = (p-1)/2 mod p How ?

jimthompson5910 (jim_thompson5910):

I don't understand what you're asking

OpenStudy (loser66):

below the line : "well I think it's enough....." you stated : (p-3)! = (p-1)/2 mod p. My question is there, how?

jimthompson5910 (jim_thompson5910):

that's what you're given

OpenStudy (loser66):

That is the statement we need to prove, not given information.

jimthompson5910 (jim_thompson5910):

I'm saying that the thing you need to prove connects to `(p-1)! =-1 (mod p)` (in the work shown above)

OpenStudy (loser66):

You meant we go back ward? A = B B leads to C and if C true then B true then A = B hold, right?

jimthompson5910 (jim_thompson5910):

what we need to prove: `(p-3)! = (p-1)/2 (mod p)` that's equivalent to saying "we need to prove `2(p-3)! = (p-1) (mod p)` is true" ----------------------------------------------- from above `(p-1)! =-1 (mod p)` turns into `2(p-3)! = -1 (mod p)` since `(p-1)! =-1 (mod p)` is true, this means `2(p-3)! = -1 (mod p)` is true so `2(p-3)! = -1 (mod p)` has been proven true leading to `(p-3)! = (p-1)/2 (mod p)` being proven true

OpenStudy (loser66):



jimthompson5910 (jim_thompson5910):